# Self-similarity on 4d cubic lattice

### v1

@@ 1 Introduction
In this paper, some algebra is developed, dealing with four-dimensional cubic lattices having a “spin” at each of its edges, and a copy of a $4\times 4$ matrix
@@ 2 General case, with indeterminates over $\bm{\mathbb{F}_{2}}$
We put a representing polynomial depending on four indeterminates $u_{j}$,  $i=1,\ldots,4$, in correspondence to a vector in each of the spaces $\mathbf{V}_{j}$ according to the following principle: for a vector $\begin{pmatrix}x_{1}&\ldots&x_{8}\end{pmatrix}\in\mathbf{V}_{j}$, take all $u_{i}$ except $u_{j}$, not changing their order, and denote them as $v_{1},v_{2},v_{3}$ (for instance, if $j=1$, then $v_{1}=u_{2}$, $v_{2}=u_{3}$, and $v_{3}=u_{4}$). The polynomial is, by definition,
$x_{1}v_{1}v_{2}v_{3}+x_{2}v_{1}v_{2}+x_{3}v_{1}v_{3}+x_{4}v_{1}+x_{5}v_{2}v_{3}+x_{6}v_{2}+x_{7}x_{3}+x_{8}.$ (7)
@@
of its intersections $\mathbf{W}_{j}=\mathbf{W}\cup\mathbf{V}_{j}$ with our thick spaces $\mathbf{V}_{j}$. Also, each $\mathbf{W}_{j}$ is of course spanned be vectors $\mathsf{e}_{i}^{(j)}$,  $i=1,\ldots,4$. We will be using a particular case of spaces $\mathbf{W}_{j}$ below in Proposition 3.

### v2

# Open Communications in Nonlinear Mathematical Physics   Special Issue: Hietarinta, 2025/2026 pp 1B

@@ 1 Introduction
This paper is significantly inspired by Hietarinta’s work [3]. We develop some algebra, dealing with four-dimensional cubic lattices having a “spin” at each of its edges, and a copy of a $4\times 4$ matrix
@@ 2 General case, with indeterminates over $\bm{\mathbb{F}_{2}}$
We put a representing polynomial depending on four indeterminates $u_{j}$,  $j=1,\ldots,4$, in correspondence to a vector in each of the spaces $\mathbf{V}_{j}$ according to the following principle: for a vector $\begin{pmatrix}x_{1}&\ldots&x_{8}\end{pmatrix}\in\mathbf{V}_{j}$, take all $u_{i}$ except $u_{j}$, not changing their order, and denote them as $v_{1},v_{2},v_{3}$ (for instance, if $j=1$, then $v_{1}=u_{2}$, $v_{2}=u_{3}$, and $v_{3}=u_{4}$). The polynomial is, by definition,
$x_{1}v_{1}v_{2}v_{3}+x_{2}v_{1}v_{2}+x_{3}v_{1}v_{3}+x_{4}v_{1}+x_{5}v_{2}v_{3}+x_{6}v_{2}+x_{7}v_{3}+x_{8}.$ (7)
@@
of its intersections $\mathbf{W}_{j}=\mathbf{W}\cap\mathbf{V}_{j}$ with our thick spaces $\mathbf{V}_{j}$. Also, each $\mathbf{W}_{j}$ is of course spanned be vectors $\mathsf{e}_{i}^{(j)}$,  $i=1,\ldots,4$. We will be using a particular case of spaces $\mathbf{W}_{j}$ below in Proposition 3.